Is R = {[a a+b
a+b b ] |a,b belongs to Z} a subring of M2(Z)? Why or why not?
1
Expert's answer
2020-04-08T11:45:35-0400

"\begin{pmatrix}
 1 & 3 \\
 3 & 2
\end{pmatrix}\in R" and "\begin{pmatrix}
 1 & 2 \\
 2 & 1
\end{pmatrix}\in R", but "\begin{pmatrix}
 1 & 3 \\
 3 & 2
\end{pmatrix}\begin{pmatrix}
 1 & 2 \\
 2 & 1
\end{pmatrix}=\begin{pmatrix}
 7 & 5 \\
 7 & 8
\end{pmatrix}\not\in R". So "R" is not a subring of "M_2(\mathbb Z)".